OpenStudy (anonymous):

Compound interest: Find the time required for an investment of $5,000 to go to $8,000 at an interest rate 7.5% per year, compounded quarterly.

OpenStudy (anonymous):

6.325284 years = 1/4 * Ln(8000/5000) / Ln(1 + 0.075/4)

OpenStudy (anonymous):

can you show using the actual compound interest formula without skipping steps please

OpenStudy (anonymous):

P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year

OpenStudy (anonymous):

yeah I understand, CAN YOU JUST SHOW ME THE FORMULA WITH THE NUMBERS IN THE CORRECT SPACES

OpenStudy (anonymous):

5000 * (1+0.075/4)^4t = 8000 so (1 + 0.075/4)^4t=8000/5000 take a log of both sides ...

OpenStudy (anonymous):

M = P( 1 + i )n

OpenStudy (anonymous):

log((1+0.075)^4t)=4t*log(1+0.075/4) so 4t=log(8000/5000)/log(1+0.075/4) and t = 1/4*log(8000/5000)/log(1+0.075/4)

OpenStudy (anonymous):

im not getting the same answer, your writing isnt very clear.